Please help me solve this equation:
1-7/x-18/x^2=0

Answer by stanbon(75887)

1-7/x-18/x^2=0
--------------------
Multiply thru by x^2 to get:
x^2-7x-18 = 0
(x-9)(x+2) = 0
x = 9, or x=-2
===================
Cheers,
Stan H.